5years from now annu will be 24years old. The current sum of the ages of annu n aarti is 49. How old is aarti right now?
Answer by stanbon(75887)

5years from now annu will be 24years old. The current sum of the ages of annu n aarti is 49. How old is aarti right now?
----
Equations:
u + 5 = 24
u + a = 49
-----------------
Substitute for "u" and solve for "a"::
19 + a = 49
a = 30 yrs ( aarti's age now )
----------
Cheers,
Stan H.
-----------